OpenStudy (anonymous):

Solve the given system of equation by substitution method. 5x+y=15 6x-2y=2

OpenStudy (raden):

from the 1st equ. 5x+y=15 ----> y = 15 - 5x, subtitute this y to the 2nd equ : 6x-2y=2 6x - 2(15 - 5x) = 2 6x - 30 + 10x = 2 16x = 30 + 2 16x = 32 x = 32/16 = 2

OpenStudy (raden):

now, subtitute the x to y = 15 - 5x, giving us y = 15 - 5(2) y = 15 - 10 y = 5

OpenStudy (raden):

therefore, the solution be (2,5)

OpenStudy (anonymous):

you can also find the third equation first which will be a long process